# A Framework for Curriculum Transformation in Quantum Information Science and Technology Education

### v2

@@ Abstract
The field of Quantum Information Science & Technology (QIST)is booming. Due to this, many new educational courses and university programs are needed in order to prepare a workforce for the developing industry. Owing to its specialist nature, teaching approaches in this field can suffer with being disconnected from the substantial degree of science education research which aims to support the best approaches to teaching in Science, Technology, Engineering & Mathematics (STEM)fields. In order to connect these two communities with a pragmatic and repeatable methodology, we have synthesised this educational research into a decision-tree based theoretical model for the transformation of QIST curricula, intended to provide a didactical perspective for practitioners. The Quantum Curriculum Transformation Framework (QCTF)consists of four steps: 1. choose a topic, 2. choose one or more targeted skills, 3. choose a learning goal and 4. choose a teaching approach that achieves this goal. We show how this can be done using an example curriculum and more specifically quantum teleportation as a basic concept of quantum communication within this curriculum. By approaching curriculum creation and transformation in this way, educational goals and outcomes are more clearly defined which is in the interest of the individual and the industry alike. The framework is intended to structure the narrative of QIST teaching, and with future testing and refinement it will form a basis for further research in the didactics of QIST.
@@ 1 Introduction
The paper is structured as follows: In Sect. 2, we summarise the basis for choosing educational content (the “what”), which makes use of the European Competence Framework (CF)for Quantum Technologies [15]. Then, three targeted skills areas are introduced in Sect. 3, as both a motivation and a means of communicating the content chosen to teach. In Sect. 4, a modified Bloom’s taxonomy is described for use in defining learning goals. In Sect. 5, Ainsworth’s Designs, Functions, Tasks (DeFT)framework is summarised and applied to the QIST context in Sect. 6. Lastly, we show a possible explicit teaching approach to quantum teleportation based on our framework in Sect. 7. Throughout this work, we will stick to examples to ground the QCTF and convey the meaning of the didactical theory in the QIST context. Quantum teleportation is chosen as a topic specifically because it makes use of various concepts central to QIST: entanglement, measurements and unitary operations in three-qubit systems and has many applications [16]. In order to make the ordering of the framework’s sections more explicit and intuitive for practitioners, we roughly structure it into a “what” - “why” - “how” format, in which we first choose content to be taught, then formulate specific teaching goals for the interaction, and then consider the specific teaching approaches which best address these goals while also respecting the diversity of student backgrounds. As such, our approach is akin to a backwards design process [17].
@@ 5.1 Functions
The Function layer within this framework describes three main roles that MER s can play: Firstly, to complement, which means to provide various perspectives and information on the same concept so that learners can benefit from the specific advantages of each representation and to support different processes, i.e. strategies and individual differences of learners. For example, the Bloch sphere provides learners with an additional strategy of reasoning about the effect of quantum gates geometrically, as rotations on the sphere [87], complementing calculating in the computational basis typical Dirac notation. Another possible function MER s is to constrain interpretation i.e. to specify ambiguous or complex information, e.g. coming from text, for example with symbols or mathematical formulas. This can be an inherent property of the representation, or can be by familiarity of the learner with the representation that is used. And lastly to construct deeper understanding, i.e. to support learners in achieving new insights on a concept which would be difficult with only one representation [84]. The Bloch sphere would be another example of this as it visualises two complex numbers, utilising the redundancy of the global phase, to enable new visual insights on single-qubit quantum systems, sometimes even leading to scientific advances, e.g. in the context of Rabi oscillations [88, 89]. In these various ways, different MER s should be chosen carefully and given as tools to support learners in achieving the levels in Bloom’s taxonomy that the educators want them to achieve. For example, in order to select the best hardware for a given quantum algorithm, one could analyse coupling maps with gate errors like the one in [90, Fig. 5], while, on the other hand, this representation is not suitable to describe the mathematics of any quantum algorithm.
@@ 5.3 Tasks
All of these effects have an influence on Extraneous cognitive load (ECL) which is the cognitive load imposed on the learner by the way the information is provided, i.e. the Design, and not the complexity of the information itself which is referred to as Intrinsic cognitive load (ICL) [102]. When ECL is reduced, learners can focus more on learning the content itself. Especially if the ICL of a given task is high, the negative effects of high ECL can become stronger (element-interactivity effect [103]). Even though ECL is usually seen as not desirable, some representations might be necessary to learn even though they are not the most efficient ECL-wise, because they are widely accepted by the academic community and broadly used for communicating scientific concepts. An example of this is the Dirac notation that should be taught in QIST courses regardless of the (possible) ECL that it brings along vs., e.g. the vector notation. In the following, the DeFT framework is applied to the QIST context in more detail.
@@ 7.1 Level of Inquiry
Research suggests that conceptual understanding and subject knowledge are best fostered in more closed, structured inquiry based tasks [124], whilst scientific thinking and inquiry skills are best fostered with more open, unstructured tasks [126]. This may be explained by the greater cognitive load associated with the more open levels of inquiry [126, 129], leaving less cognitive ability to develop conceptual understanding or commit ideas to long-term memory. For this reason, educators must choose carefully the level of inquiry in order to match their learning goals. In the field of QIST, there have been many recent examples of interactive tools which create possibilities for teaching approaches which span the four levels of inquiry [130]. For example, the visual and gamified simulation tool for Quantum Computing, Quantum Odyssey [131], is designed with a coherent structure of stand-alone scenarios addressing individual concepts, such as particular gate operations and communication protocols. These can be used to create more structured and confirmatory inquiry scenarios in which participants are both introduced to concepts and walked through them towards attaining particular targeted knowledge and understanding around the concepts. On the other hand, the Quantum Composer tool is suited to truly open inquiry as it provides a blank slate interface in which staff and students may construct any one-dimensional quantum system and investigate its properties. In fact the tool has been used to support research in real problems in QIST [132, 133], which demonstrates how this form of inquiry can support scientific thinking as a learning outcome. Both of these tools provide a means of visual representation of concepts in QIST, but it should be noted that IBL models need not only use visualisation, but other forms of representation as described in Sect. 5.
@@ 7.3 Scaffolding
Especially in the QIST context, we argue that various scaffolding techniques could be beneficial, because they consider individual academic backgrounds by definition. In addition, the concepts in QIST are quite abstract while the methods to deal with these concepts are often new to students, with possibly discouraging effect. Lastly, the field of QIST is new and therefore, knowledge acquisition often requires students to do their own research, reading scientific research articles that are not designed with them as a target audience in mind. Scaffolding could, e.g. be provided in the form of providing on demand suitable visualisations, encouraging peer-to-peer instruction and cooperativity (as covered in the previous section), providing help with representations and how to perform calculations, concept maps (e.g. connecting content with content from different targeted skill areas or different areas of the CF) aiding in research tasks by providing strategies and foundational material for understanding or close interactions with researchers (e.g. interviews).
@@ 7.4 Teaching Approach to Quantum Teleportation
Step
Skill Area
Up to this level in Bloom’s taxonomy
Activity
Theory
Understand
Listening to short, conceptual presentation on quantum teleportation and discussion
Analytics
Apply
Calculate necessary gates, choose proper representations
Experiment
Apply
Plan an experiment and sketch an experimental setup
Real World Application
Evaluate
Discuss real world application of quantum teleportation in quantum networks in a presentation
Computation & Simulation or Experiment
Remember
Carry out experiments or outlook on further use cases of quantum teleportation in quantum computing & programming the algorithm
@@ 8 Outlook and Discussion
We emphasise that the framework is a guiding structure, intended to distil the results of educational research into a pragmatic approach for educators. Whilst most effectively applied in full, educators may also pick aspects from the framework to address individually - for example refining primarily their level of inquiry, or choosing topics using the CF. The framework is also open to refinement and iteration based on community feedback, as more and more educational material is developed in this field, and in particular with an eye to a didactically-sound approach. One particular refinement which may be beneficial in the future is the inclusion of proficiency levels in the “what” stage, while choosing content. Suggested proficiency levels are currently included in the Competence Framework [18], providing a rough equivalency between content difficulty and qualification level. However, as yet they are not based on educational literature and thus were not included in the framework in this iteration. The refinement of the proficiency levels, as the framework itself, we recommend for future work in the course of the flagship European Quantum Technology education projects, DigiQ [173] and QTIndu [174].

### v3

@@ Abstract
The field of Quantum Information Science & Technology (QIST)is booming. Due to this, many new educational courses and university programs are needed in order to prepare a workforce for the developing industry. Owing to its specialist nature, teaching approaches in this field can easily become disconnected from the substantial degree of science education research which aims to support the best approaches to teaching in Science, Technology, Engineering & Mathematics (STEM)fields. In order to connect these two communities with a pragmatic and repeatable methodology, we have synthesised this educational research into a decision-tree based theoretical model for the transformation of QIST curricula, intended to provide a didactical perspective for practitioners. The Quantum Curriculum Transformation Framework (QCTF)consists of four steps: 1. choose a topic, 2. choose one or more targeted skills, 3. choose a learning goal and 4. choose a teaching approach that achieves this goal. We show how this can be done using an example curriculum and more specifically quantum teleportation as a basic concept of quantum communication within this curriculum. By approaching curriculum creation and transformation in this way, educational goals and outcomes are more clearly defined which is in the interest of the individual and the industry alike. The framework is intended to structure the narrative of QIST teaching, and with future testing and refinement it will form a basis for further research in the didactics of QIST.
@@ 1 Introduction
The paper is structured as follows: In Sect. 2, we summarise the basis for choosing educational content (the “what”), which makes use of the European Competence Framework (CF)for Quantum Technologies [15, 16]. Then, three targeted skills areas are introduced in Sect. 3, as both a motivation and a means of communicating the content chosen to teach. In Sect. 4, a modified Bloom’s taxonomy is described for use in defining learning goals. In Sect. 5, Ainsworth’s Designs, Functions, Tasks (DeFT)framework is summarised and applied to the QIST context in Sect. 6. Lastly, we show a possible explicit teaching approach to quantum teleportation based on our framework in Sect. 7. Throughout this work, we will stick to examples to ground the QCTF and convey the meaning of the didactical theory in the QIST context. Quantum teleportation is chosen as a topic specifically because it makes use of various concepts central to QIST: entanglement, measurements and unitary operations in three-qubit systems and has many applications [17]. In order to make the ordering of the framework’s sections more explicit and intuitive for practitioners, we roughly structure it into a “what” - “why” - “how” format, in which we first choose content to be taught, then formulate specific teaching goals for the interaction, and then consider the specific teaching approaches which best address these goals while also respecting the diversity of student backgrounds. As such, our approach is akin to a backwards design process [18].
@@ 5.1 Functions
The Function layer within this framework describes three main roles that MER s can play: Firstly, to complement, which means to provide various perspectives and information on the same concept so that learners can benefit from the specific advantages of each representation and to support different processes, i.e. strategies and individual differences of learners. For example, the Bloch sphere provides learners with an additional strategy of reasoning about the effect of quantum gates geometrically, as rotations on the sphere [88], complementing calculating in the computational basis typical Dirac notation. Another possible function MER s is to constrain interpretation i.e. to specify ambiguous or complex information, e.g. coming from text, for example with symbols or mathematical formulas. This can be an inherent property of the representation, or can be by familiarity of the learner with the representation that is used. And lastly to construct deeper understanding, i.e. to support learners in achieving new insights on a concept which would be difficult with only one representation [85]. The Bloch sphere would be another example of this as it visualises two complex numbers, utilising the redundancy of the global phase, to enable new visual insights on single-qubit quantum systems, sometimes even leading to scientific advances, e.g. in the context of Rabi oscillations [89, 90]. In these various ways, different MER s should be chosen carefully and given as tools to support learners in achieving the levels in Bloom’s taxonomy that the educators want them to achieve. For example, in order to select the best hardware for a given quantum algorithm, one could analyse coupling maps with gate errors like the one in [91, Fig. 5], while, on the other hand, this representation is not suitable to describe the mathematics of any quantum algorithm.
@@ 5.3 Tasks
All of these effects have an influence on Extraneous cognitive load (ECL) which is the cognitive load imposed on the learner by the way the information is provided, i.e. the Design, and not the complexity of the information itself which is referred to as Intrinsic cognitive load (ICL) [103]. When ECL is reduced, learners can focus more on learning the content itself. Especially if the ICL of a given task is high, the negative effects of high ECL can become stronger (element-interactivity effect [104]). Even though ECL is usually seen as not desirable, some representations might be necessary to learn even though they are not the most efficient ECL-wise, because they are widely accepted by the academic community and broadly used for communicating scientific concepts, an example of this being the Dirac notation. In the following, the DeFT framework is applied to the QIST context in more detail.
@@ 7.1 Level of Inquiry
Research suggests that conceptual understanding and subject knowledge are best fostered in more closed, structured inquiry based tasks [125], whilst scientific thinking and inquiry skills are best fostered with more open, unstructured tasks [127]. This may be explained by the greater cognitive load associated with the more open levels of inquiry [127, 130], leaving less cognitive ability to develop conceptual understanding or commit ideas to long-term memory. For this reason, educators must choose carefully the level of inquiry in order to match their learning goals. In the field of QIST, there have been many recent examples of interactive tools which create possibilities for teaching approaches which span the four levels of inquiry [131]. For example, the visual and gamified simulation tool for Quantum Computing, Quantum Odyssey [132], is designed with a coherent structure of stand-alone scenarios addressing individual concepts, such as particular gate operations and communication protocols. These can be used to create more structured and confirmatory inquiry scenarios in which participants are both introduced to concepts and walked through them towards attaining particular targeted knowledge and understanding around the concepts. In addition, Odyssey has an editor mode in which users can create quantum algorithms from scratch using custom gates, therefore allowing for more open inquiry. The Quantum Composer tool is also effective for open inquiry, as it provides a blank slate interface in which staff and students may construct any one-dimensional quantum system and investigate its properties. In fact, the tool has been used to support research in real problems in QIST [133, 134], which demonstrates how this form of inquiry can support scientific thinking as a learning outcome. Both of these tools provide a means of visual representation of concepts in QIST, but it should be noted that IBL models need not only use visualisation, but other forms of representation as described in Sect. 5.
@@ 7.3 Scaffolding
Especially in the QIST context, we argue that various scaffolding techniques could be beneficial, because they consider individual academic backgrounds by definition. In addition, the concepts in QIST are quite abstract while the methods to deal with these concepts are often new to students, with possibly discouraging effect. Lastly, the field of QIST is new and therefore, knowledge acquisition often requires students to do their own research, reading scientific research articles that are not designed with them as a target audience in mind. Scaffolding could, e.g. be provided in the form of providing on demand suitable visualisations, encouraging peer-to-peer instruction and cooperativity (as covered in the previous section), providing help with representations and how to perform calculations, concept maps (e.g. connecting content with content from different targeted skill areas or different domains of the CF) aiding in research tasks by providing strategies and foundational material for understanding or close interactions with researchers (e.g. interviews).
@@ 7.4 Teaching Approach to Quantum Teleportation
Step Skill Area Up to this level in Bloom’s taxonomy Activity
 (i) Theory Understand Listening to short, conceptual presentation on quantum teleportation and discussion
(ii) Analytics Apply Calculate necessary gates, choose proper representations
(iii) Experiment Apply Plan an experiment and sketch an experimental setup
(iv) Real World Application Evaluate Discuss real world application of quantum teleportation in quantum networks in a presentation
(v) Computation & Simulation or Experiment Remember Carry out experiments or outlook on further use cases of quantum teleportation in quantum computing & programming the algorithm
@@ 8 Outlook and Discussion
We emphasise that the framework is a guiding structure, intended to distil the results of educational research into a pragmatic approach for educators. Whilst most effectively applied in full, educators may also pick aspects from the framework to address individually - for example refining primarily their level of inquiry, or choosing topics using the CF. The framework is also open to refinement and iteration based on community feedback, as more and more educational material is developed in this field, and in particular with an eye to a didactically-sound approach. One particular refinement which may be beneficial in the future is the inclusion of proficiency levels in the “what” stage, while choosing content. Suggested proficiency levels were included in the Competence Framework version 2.0 [19], providing a rough equivalency between content difficulty and qualification level. Since then, the framework was refined on the basis of interviews with industry representatives, adding 9 qualification profiles that could further guide development of curricula [16]. The refinement of the proficiency levels, as the framework itself, we recommend for future work in the course of the flagship European Quantum Technology education projects, DigiQ [174] and QTIndu [175].
@@ Acknowledgements
We thank Franziska Greinert from the TU Braunschweig for valuable discussion.